At a break-even point of 300 units, variable costs were $480 and fixed costs were $195. What will the 301st unit sold contribute to operating profits before income taxes?
Answer #1

Variable cost per unit = 480/300 = 1.60

Breakeven point = Fixed cost / Contribution margin per unit

195/(Sales price - 1.60) = 300

Sales price = 2.25

Contribution per unit = 2.25 - 1.60

= 0.65